# The Nikodym property and filters on $\omega$

### v1

@@ 1. Introduction
    there is a density submeasure $\varphi$ on $\omega$ such that $\varphi(\omega)=\infty$ and $*F\subseteq\operatorname{Exh}(\varphi)^{*}$;
@@
    there is a non-pathological lsc submeasure $\varphi$ on $\omega$ such that $\varphi(\omega)=\infty$ and $*F\subseteq\operatorname{Exh}(\varphi)^{*}$.
@@ 3. Anti-Nikodym sequences of measures on spaces $N_{F}$
Proposition 3.2.
  1. (1)
    $\lim_{n\to\infty}\|\mu_{n}\|=\infty$,
  2. (2)
    $\lim_{n\to\infty}\mu_{n}(N_{F})=0$,
  3. (3)
Proof.
Assume that $\langle\mu_{n}\colon n\in\omega\rangle$ is an AN-sequence on $N_{F}$. Then, $(1)$ and $(2)$ hold by the definition. Assume, for the sake of contradiction, that for some $A\in F$ there is a subsequence $\langle\mu_{n_{k}}:\ k\in\omega\rangle$ such that $\big{\|}\mu_{n_{k}}\restriction(\omega\mathbin{\backslash}A)\big{\|}>0$ for every $k\in\omega$ and $\lim_{k\to\infty}\big{\|}\mu_{n_{k}}\restriction(\omega\mathbin{\backslash}A)\big{\|}>0$. Then, proceeding as in the proof of Proposition 6.1, one can prove that the sequence
$\big{\langle}\big{(}\mu_{n_{k}}\restriction(\omega\mathbin{\backslash}A)\big{)}/\|\mu_{n_{k}}\restriction(\omega\mathbin{\backslash}A)\|\colon k\in\omega\big{\rangle}$
is a BJN-sequence on the discrete clopen subspace $\omega\mathbin{\backslash}A$ of $N_{F}$, contradicting [22, Lemma 5.1.(4)]. This proves $(3)$.
@@
We first define inductively a strictly increasing sequence $\langle n_{k}:\ k\in\omega\rangle$ of natural numbers and an increasing sequence $\langle A_{k}:\ k\in\omega\rangle$ of subsets of $N_{F}$ satisfying $A_{k}=\big{[}0,\max\big{(}\bigcup_{i=0}^{k-1}\operatorname{supp}(\mu_{n_{i}})\cap\omega\big{)}\big{]}$ and $|\mu_{n_{k}}|(A_{k})<1/k$ for every $k\in\omega$. We start with $n_{0}=0$ and $A_{0}=\emptyset$. Assume now that for some $k\in\omega_{+}$ we have defined natural numbers $n_{0}<\ldots<n_{k-1}$ and subsets $A_{0}\subseteq\ldots\subseteq A_{k-1}$ of $N_{F}$ as required. The set $A_{k}=\big{[}0,\max\big{(}\bigcup_{i=0}^{k-1}\operatorname{supp}(\mu_{n_{i}})\cap\omega\big{)}\big{]}$ is a finite subset of $\omega$, so $A_{k}\in\operatorname{Clopen}(N_{F})$. As $\langle\mu_{n}\colon n\in\omega\rangle$ is an AN-sequence, there exists $n_{k}>n_{k-1}$ such that $|\mu_{n_{k}}|(A_{k})<1/k$.
@@ 4. The Nikodym property and submeasures on $\omega$
Then $\varphi$ is a density submeasure. We have $\varphi(\omega)=\sup_{n\in\omega}\mu_{n}(\omega)=\infty$ by condition $(2)$ from Theorem 1.3. It is fairly easy to see that $*F\subseteq\operatorname{Exh}(\varphi)^{*}$ (cf. [22, Theorem 6.2]).
@@
$(3)\Rightarrow(1)$ Assume that there is a non-pathological lsc submeasure $\varphi$ such that $\varphi(\omega)=\infty$ and $*F\subseteq\operatorname{Exh}(\varphi)^{*}$. By Remark 3.1, it is enough to prove that $*\operatorname{Exh}(\varphi)^{*}$ does not have the Nikodym property.
@@
We claim that the sequence $\langle\mu_{k}:\ k\in\omega\rangle$ satisfies the equality $\lim_{k\to\infty}\mu_{k}(A)=0$ for every $A\in\operatorname{Exh}(\varphi)$, and thus, by Theorem 1.3, the filter $*\operatorname{Exh}(\varphi)^{*}$ does not have the Nikodym property. Let $A\in\operatorname{Exh}(\varphi)$ and $\varepsilon>0$. Since $\lim_{n\to\infty}\varphi\big{(}A\mathbin{\backslash}[0,n]\big{)}=0$, there is $M\in\omega$ such that $\varphi\big{(}A\mathbin{\backslash}[0,M]\big{)}<\varepsilon$. Let $k\in\omega$ be such that $n_{k}>M$. It holds:
@@ 5. The Nikodym property and the Katětov order
    there is a density submeasure $\varphi$ on $\omega$ such that $\varphi(\omega)=\infty$ and $*F\leq_{K}\operatorname{Exh}(\varphi)^{*}$;
@@
    there is a non-pathological lsc submeasure $\varphi$ on $\omega$ such that $\varphi(\omega)=\infty$ and $*F\leq_{K}\operatorname{Exh}(\varphi)^{*}$.
@@
$(1)\Rightarrow(2)$ Let $F$ be a filter on $\omega$ without the Nikodym property. By Theorem 1.3, there is a density submeasure $\varphi$ on $\omega$ such that $\varphi(\omega)=\infty$ and $*F\subseteq\operatorname{Exh}(\varphi)^{*}$, and so $*F\leq_{K}\operatorname{Exh}(\varphi)^{*}$.
@@
$(3)\Rightarrow(1)$ Let $\varphi$ be a non-pathological lsc submeasure on $\omega$ such that $\varphi(\omega)=\infty$ and $*F\leq_{K}\operatorname{Exh}(\varphi)^{*}$. By Theorem 1.3, the filter $*\operatorname{Exh}(\varphi)^{*}$ does not have the Nikodym property, thus by Proposition 5.2 the filter $F$ does not have the Nikodym property. ∎
@@
Note that the assumption that ${\mathcal{I}}$ is a density ideal was only used in the proof of implication $(7)\Rightarrow(3)$ of Theorem 1.3.
@@ 5.1. Cofinal structure of $(\mathcal{A}{\mathcal{N}},\leq_{K})$
$(1)$ Let ${\mathcal{I}}\in\mathcal{A}{\mathcal{N}}$. By Theorem 1.3, there is a density ideal $\operatorname{Exh}\big{(}\sup_{n\in\omega}\lambda_{n}\big{)}$ for some sequence $\langle\lambda_{n}\colon n\in\omega\rangle$ of finitely supported measures with disjoint supports such that $\|\lambda_{n}\|\to\infty$ and ${\mathcal{I}}\subseteq\operatorname{Exh}\big{(}\sup_{n\in\omega}\lambda_{n}\big{)}$. Without losing of generality we may assume that $\|\lambda_{n}\|\geq n+1$ for all $n\in\omega$, as by taking a subsequence of the sequence $\langle\lambda_{n}\colon n\in\omega\rangle$ we can only enlarge the ideal $\operatorname{Exh}\big{(}\sup_{n\in\omega}\lambda_{n}\big{)}$.
@@
$(2)$ Let $g,h\in\omega^{\omega}$ be such that $2n^{2}\cdot g(n)\leq h(n)$ for all but finitely many $n\in\omega$. It follows immediately from Lemma 5.8 that $\Phi(g)\leq_{K}\Phi(h)$. ∎
@@
If ${\mathcal{F}}$ is a dominating family in $\omega^{\omega}$, then Φ[]normal-$\Phi[{\mathcal{F}}]$ is $\leq_{K}$-dominating in $\mathcal{A}{\mathcal{N}}$, i.e. for any ${\mathcal{I}}\in\mathcal{A}{\mathcal{N}}$ there is $f\in{\mathcal{F}}$ such that ${\mathcal{I}}\leq_{K}\Phi(f)$.
@@ 6. $\mathcal{A}{\mathcal{N}}$ vs ${\mathcal{B}}{\mathcal{J}}{\mathcal{N}}{\mathcal{P}}$
We will now summarize differences between the classes of ideals $\mathcal{A}{\mathcal{N}}$ and ${\mathcal{B}}{\mathcal{J}}{\mathcal{N}}{\mathcal{P}}$. As we said in Introduction, these are connected with differences between the Nikodym property and the Grothendieck property of Boolean algebras. We start with the following result, stating that $\mathcal{A}{\mathcal{N}}$ is a subclass of ${\mathcal{B}}{\mathcal{J}}{\mathcal{N}}{\mathcal{P}}$.
@@
$\sum_{i=1}^{m}|\alpha_{i}|\cdot|\mu_{n}(A_{i})|<\varepsilon/2.$
@@
$\leq\Big{\|}f-\sum_{i=1}^{m}\alpha_{i}\chi_{A_{i}}\Big{\|}_{\infty}\cdot\|\mu_{n}\|+\sum_{i=1}^{m}|\alpha_{i}|\cdot|\mu_{n}(A_{i})|<\varepsilon/2+\varepsilon/2=\varepsilon,$
@@
    there is a density submeasure $\varphi$ such that $*F\subseteq\operatorname{Exh}(\varphi)^{*}$,
@@
    there is a non-pathological lsc submeasure $\varphi$ such that $*F\subseteq\operatorname{Exh}(\varphi)^{*}$,
@@ 9.1. Non-totally bounded non-pathological ideals
Does there exist a non-pathological lsc submeasure $\varphi$ such that the ideal $\operatorname{Exh}(\varphi)$ is not totally bounded and $*\operatorname{Exh}(\varphi)^{*}$ has the Nikodym property?

### v2

@@ 1. Introduction
    there is a density submeasure $\varphi$ on $\omega$ such that $\varphi(\omega)=\infty$ and $F\subseteq\operatorname{Exh}(\varphi)^{*}$;
@@
    there is a non-pathological lsc submeasure $\varphi$ on $\omega$ such that $\varphi(\omega)=\infty$ and $F\subseteq\operatorname{Exh}(\varphi)^{*}$.
@@ 3. Anti-Nikodym sequences of measures on spaces $N_{F}$
We first define inductively a strictly increasing sequence $\langle n_{k}:\ k\in\omega\rangle$ of natural numbers and an increasing sequence $\langle A_{k}:\ k\in\omega\rangle$ of subsets of $N_{F}$ satisfying $A_{k}=\big{[}0,\max\big{(}\bigcup_{i=0}^{k-1}\operatorname{supp}(\mu_{n_{i}})\cap\omega\big{)}\big{]}$ and $|\mu_{n_{k}}|(A_{k})<1/k$ for every $k\in\omega$. We start with $n_{0}=0$ and $A_{0}=\emptyset$. Assume now that for some $k\in\omega_{+}$ we have defined natural numbers $n_{0}<\ldots<n_{k-1}$ and subsets $A_{0}\subseteq\ldots\subseteq A_{k-1}$ of $N_{F}$ as required. The set $A_{k}=\big{[}0,\max\big{(}\bigcup_{i=0}^{k-1}\operatorname{supp}(\mu_{n_{i}})\cap\omega\big{)}\big{]}$ is a finite subset of $\omega$. As $\langle\mu_{n}\colon n\in\omega\rangle$ is an AN-sequence and singletons are clopen, there exists $n_{k}>n_{k-1}$ such that $|\mu_{n_{k}}|(A_{k})<1/k$.
@@
Proposition 3.3.
  1. (1)
    $\lim_{n\to\infty}\|\mu_{n}\|=\infty$,
  2. (2)
    $\lim_{n\to\infty}\mu_{n}(N_{F})=0$,
  3. (3)
Proof.
Let $\langle\mu_{n}\colon n\in\omega\rangle$ be an AN-sequence on $N_{F}$. Then, $(1)$ and $(2)$ hold by the definition. By Proposition 3.2, without losing of generality we may assume that the sequence $\langle\mu_{n}\colon n\in\omega\rangle$ is disjointly supported. Assume, for the sake of contradiction, that for some $A\in F$ there is a subsequence $\langle\mu_{n_{k}}:\ k\in\omega\rangle$ such that $\lim_{k\to\infty}\big{\|}\mu_{n_{k}}\restriction(\omega\mathbin{\backslash}A)\big{\|}=C$ for some $C>0$. We may additionally assume that $\big{\|}\mu_{n_{k}}\restriction(\omega\mathbin{\backslash}A)\big{\|}\geq C/2$ for every $k\in\omega$. For each $k\in\omega$ there exists $F_{k}\subseteq\operatorname{supp}\big{(}\mu_{n_{k}}\restriction(\omega\mathbin{\backslash}A)\big{)}$ satisfying $|\mu_{n_{k}}(F_{k})|\geq C/4$. Since $\omega\mathbin{\backslash}A$ is a discrete clopen subspace of $N_{F}$, the set $F=\bigcup_{k\in\omega}F_{k}$ is clopen in $N_{F}$. But $\lim_{k\to\infty}|\mu_{n_{k}}(F)|\geq C/4$, contradicting the fact that $\langle\mu_{n_{k}}:\ k\in\omega\rangle$ is an AN-sequence on $N_{F}$.
@@ 4. The Nikodym property and submeasures on $\omega$
Then $\varphi$ is a density submeasure. We have $\varphi(\omega)=\sup_{n\in\omega}\mu_{n}(\omega)=\infty$ by condition $(2)$ from Theorem 1.3. It is fairly easy to see that $F\subseteq\operatorname{Exh}(\varphi)^{*}$ (cf. [22, Theorem 6.2]).
@@
$(3)\Rightarrow(1)$ Assume that there is a non-pathological lsc submeasure $\varphi$ such that $\varphi(\omega)=\infty$ and $F\subseteq\operatorname{Exh}(\varphi)^{*}$. By Remark 3.1, it is enough to prove that $\operatorname{Exh}(\varphi)^{*}$ does not have the Nikodym property.
@@
We claim that the sequence $\langle\mu_{k}:\ k\in\omega\rangle$ satisfies the equality $\lim_{k\to\infty}\mu_{k}(A)=0$ for every $A\in\operatorname{Exh}(\varphi)$, and thus, by Theorem 1.3, the filter $\operatorname{Exh}(\varphi)^{*}$ does not have the Nikodym property. Let $A\in\operatorname{Exh}(\varphi)$ and $\varepsilon>0$. Since $\lim_{n\to\infty}\varphi\big{(}A\mathbin{\backslash}[0,n]\big{)}=0$, there is $M\in\omega$ such that $\varphi\big{(}A\mathbin{\backslash}[0,M]\big{)}<\varepsilon$. Let $k\in\omega$ be such that $n_{k}>M$. It holds:
@@ 5. The Nikodym property and the Katětov order
    there is a density submeasure $\varphi$ on $\omega$ such that $\varphi(\omega)=\infty$ and $F\leq_{K}\operatorname{Exh}(\varphi)^{*}$;
@@
    there is a non-pathological lsc submeasure $\varphi$ on $\omega$ such that $\varphi(\omega)=\infty$ and $F\leq_{K}\operatorname{Exh}(\varphi)^{*}$.
@@
$(1)\Rightarrow(2)$ Let $F$ be a filter on $\omega$ without the Nikodym property. By Theorem 1.3, there is a density submeasure $\varphi$ on $\omega$ such that $\varphi(\omega)=\infty$ and $F\subseteq\operatorname{Exh}(\varphi)^{*}$, and so $F\leq_{K}\operatorname{Exh}(\varphi)^{*}$.
@@
$(3)\Rightarrow(1)$ Let $\varphi$ be a non-pathological lsc submeasure on $\omega$ such that $\varphi(\omega)=\infty$ and $F\leq_{K}\operatorname{Exh}(\varphi)^{*}$. By Theorem 1.3, the filter $\operatorname{Exh}(\varphi)^{*}$ does not have the Nikodym property, thus by Proposition 5.2 the filter $F$ does not have the Nikodym property. ∎
@@ 5.1. Cofinal structure of $(\mathcal{A}{\mathcal{N}},\leq_{K})$
$(1)$. Let ${\mathcal{I}}\in\mathcal{A}{\mathcal{N}}$. By Theorem 1.3, there is a density ideal $\operatorname{Exh}\big{(}\sup_{n\in\omega}\lambda_{n}\big{)}$ for some sequence $\langle\lambda_{n}\colon n\in\omega\rangle$ of finitely supported measures with disjoint supports such that $\|\lambda_{n}\|\to\infty$ and ${\mathcal{I}}\subseteq\operatorname{Exh}\big{(}\sup_{n\in\omega}\lambda_{n}\big{)}$. Without losing of generality we may assume that $\|\lambda_{n}\|\geq n+1$ for all $n\in\omega$, as by taking a subsequence of the sequence $\langle\lambda_{n}\colon n\in\omega\rangle$ we can only enlarge the ideal $\operatorname{Exh}\big{(}\sup_{n\in\omega}\lambda_{n}\big{)}$.
@@
$(2)$. Let $g,h\in\omega^{\omega}$ be such that $2n^{2}\cdot g(n)\leq h(n)$ for all but finitely many $n\in\omega$. It follows immediately from Lemma 5.8 that $\Phi(g)\leq_{K}\Phi(h)$. ∎
@@
If ${\mathcal{F}}$ is a dominating family in $\omega^{\omega}$, then $\Phi[{\mathcal{F}}]$ is $\leq_{K}$-dominating in $\mathcal{A}{\mathcal{N}}$, i.e. for any ${\mathcal{I}}\in\mathcal{A}{\mathcal{N}}$ there is $f\in{\mathcal{F}}$ such that ${\mathcal{I}}\leq_{K}\Phi(f)$.
@@ 6. $\mathcal{A}{\mathcal{N}}$ vs ${\mathcal{B}}{\mathcal{J}}{\mathcal{N}}{\mathcal{P}}$
We will now summarize differences between the classes of ideals $\mathcal{A}{\mathcal{N}}$ and ${\mathcal{B}}{\mathcal{J}}{\mathcal{N}}{\mathcal{P}}$. As we said in Introduction, these are connected with differences between the Nikodym property and the Grothendieck property of Boolean algebras. We start with the following result, stating that $\mathcal{A}{\mathcal{N}}$ is a subclass of ${\mathcal{B}}{\mathcal{J}}{\mathcal{N}}{\mathcal{P}}$. Note that it actually follows from Theorems 1.3 and 6.2, but for the convenience of the reader and the self-containment of the paper we provide here a direct (and more analytic) argument.
@@
$\sum_{i=1}^{m}|\alpha_{i}|\cdot|\overline{\lambda_{n}}(A_{i})|<\varepsilon/2.$
@@
$\leq\Big{\|}f-\sum_{i=1}^{m}\alpha_{i}\chi_{A_{i}}\Big{\|}_{\infty}\cdot\|\overline{\lambda_{n}}\|+\sum_{i=1}^{m}|\alpha_{i}|\cdot|\overline{\lambda_{n}}(A_{i})|<\varepsilon/2+\varepsilon/2=\varepsilon,$
@@
    there is a density submeasure $\varphi$ such that $F\subseteq\operatorname{Exh}(\varphi)^{*}$,
@@
    there is a non-pathological lsc submeasure $\varphi$ such that $F\subseteq\operatorname{Exh}(\varphi)^{*}$,
@@ 9.1. Non-totally bounded non-pathological ideals
Does there exist a non-pathological lsc submeasure $\varphi$ such that the ideal $\operatorname{Exh}(\varphi)$ is not totally bounded and $\operatorname{Exh}(\varphi)^{*}$ has the Nikodym property?